The U.S. Department of Transportation maintains statistics for mishandled bags per 1,000 airline passengers. In February, XYZ airlines mishandled 4.05 bags per 1,000 passengers.
Assume Poisson distribution.

  1. What is the probability that in the next 1,000 passengers, XYZ will have at least five mishandled bags?
    1. 0.299
    2. 0.381
    3. 0.415
    4. 0.687
Answer #1

Let X denotes the number of mishandles bags in a randomly selected group of 1000 passengers.

X ~ Poisson(4.05)

The probability mass function of X is

P(X = z) = e-40-4.05- =0.1.2.

The probability that in the next 1,000 passengers, XYZ will have at least 5 mishandled bags

=P(Xgeq 5) = 1 - P(X< 5) = 1 - P(Xleq 4)

44.05 1 --「e-4.05 r!

= 1-0.619070

0.38093

ans-> B) 0.381